# Tractable Metric Spaces and the Continuity of Magnitude

### v1

@@ 3 Tractable Spaces
For example, recall that $\ell_{1}$ isometrically embeds into $L_{1}$. For $t\in\mathbb{R}_{>0}$ let $K_{t}$ be the closed convex hull of the set $\left\{{0}\right\}\cup\left\{{\tfrac{t}{2^{n}}e_{n}}\;\middle|\;{n\in\mathbb{N}_{\geq 1}}\right\}$ in $\ell_{1}$. By discussion in [8, Section 2], $K_{t}$ is compact and has finite magnitude; specifically, by [8, Remark 2.2], we have $\mathsf{mag}_{\ell_{1}}(K_{t})\leq e^{t/2}$. Consider $K_{1}$; by Proposition 3.6, it is tractable, but is clearly not contained in a finite-dimensional vector subspace. Since $t\cdot K_{1}=K_{t}$ and $\lim_{t\to 0}e^{t/2}=1$, we see that nevertheless $K_{1}$ has the one-point property.
This raises the question of whether every non-empty compact subspace of a tractable subspace of $L_{1}$ (equivalently, by Proposition 3.4: every compact tractable subspace of $L_{1}$) has the one-point property. We do not yet know the answer, and the question is related to Conjecture 5.1 below. We do note, however, that the same idea as in the proof of [8, Theorem 3.1] does not work, as there it was crucial that the compact subspace was contained in a polytope (i.e. convex hull of finitely many points, and therefore necessarily contained in a finite-dimensional vector subspace).
To wrap up this section, we note that in practice, it is not equally difficult to (dis)prove (uniform) continuity of magnitude for finite subspaces and compact subspaces. Specifically, it is usually easier to prove continuity for finite subspaces, and easier to disprove continuity for compact subspace. For tractable spaces, the following proposition gives us the connection between the two, potentially allowing us to derive the more difficult case from the easier one. We will use this in the next section.
@@
    Conversely, suppose that $\mathsf{mag}_{\mathrm{Fin}_{+}(M)}\colon\mathrm{Fin}_{+}(M)\to\mathbb{R}$ is uniformly continuous/$C$-Lipschitz. Then it has a unique continuous extension $\mathrm{Cmp}_{+}(M)\to\mathbb{R}$ which is also uniformly continuous/$C$-Lipschitz. But for positive definite spaces this extension has to be the magnitude $\mathsf{mag}_{\mathrm{Cmp}_{+}(M)}\colon\mathrm{Cmp}_{+}(M)\to\mathbb{R}$ [10].
@@ 4 Continuity of Magnitude on $\mathbb{R}$ and subspaces of $\mathbb{R}$
The goal of this section is to deduce that magnitude $\mathsf{mag}_{\mathrm{Cmp}_{+}(\mathbb{R})}\colon\mathrm{Cmp}_{+}(\mathbb{R})\to\mathbb{R}$ is continuous from the theory developed in the previous section. This is difficult to prove directly, even though the result has appeared in the literature previously, see [5]. We also prove that the restrictions of $\mathsf{mag}_{\mathrm{Cmp}_{+}(\mathbb{R})}$ to bounded subspaces are Lipschitz.
@@
The space $\mathbb{R}_{\geq 0}$ is tractable, but the magnitude maps $\mathsf{mag}_{\mathrm{Fin}_{+}(\mathbb{R}_{\geq 0})}\colon\mathrm{Fin}_{+}(\mathbb{R}_{\geq 0})\to\mathbb{R}$ and $\mathsf{mag}_{\mathrm{Cmp}_{+}(\mathbb{R}_{\geq 0})}\colon\mathrm{Cmp}_{+}(\mathbb{R}_{\geq 0})\to\mathbb{R}$ are not uniformly continuous (therefore also not Lipschitz continuous). Hence, the same is also true for every tractable metric space containing an isometric copy of $\mathbb{R}_{\geq 0}$.
@@
$=\big{(}1+2N\tanh(\tfrac{1}{2})\big{)}-\big{(}1+N\tanh(1)\big{)}=N\cdot\big{(}2\tanh(\tfrac{1}{2})-\tanh(1)\big{)}>C.$
@@
For all $a,b\in\mathbb{R}$, $a<b$, the magnitude maps $\mathsf{mag}_{\mathrm{Fin}_{+}({\mathbb{R}}_{[a,b]})}\colon\mathrm{Fin}_{+}({\mathbb{R}}_{[a,b]})\to\mathbb{R}$ and $\mathsf{mag}_{\mathrm{Cmp}_{+}({\mathbb{R}}_{[a,b]})}\colon\mathrm{Cmp}_{+}({\mathbb{R}}_{[a,b]})\to\mathbb{R}$ are Lipschitz with coefficient $1+\tfrac{b-a}{2}$.
@@
$=r\cdot\Big{(}n-\sum_{i\in I}\frac{\tanh(\tfrac{a_{i+1}-a_{i}}{2}-r)-\tanh(\tfrac{a_{i+1}-a_{i}}{2})}{-r}-\sum_{j\in J}\big{(}1-\tfrac{1}{r}\big{(}\tfrac{a_{j+1}-a_{j}}{2}-\tanh(\tfrac{a_{j+1}-a_{j}}{2})\big{)}\big{)}\Big{)}.$
@@
$=1+\sum_{k\in{\mathbb{N}}_{[1,n)}}\tanh^{2}(\tfrac{a_{k+1}-a_{k}}{2})\leq 1+\sum_{k\in{\mathbb{N}}_{[1,n)}}\tanh(\tfrac{a_{k+1}-a_{k}}{2})\leq$
@@ 5 Conclusion
Although we have shown here continuity of magnitude only for the subspaces of $\mathbb{R}$, we suspect that it holds much more generally. We state this as the following conjecture, which is a generalization of [3, Conjecture 1.3].

### v2

@@ 3 Tractable Spaces
For example, recall that $\ell_{1}$ isometrically embeds into $L_{1}$. For $t\in\mathbb{R}_{>0}$ let $K_{t}$ be the closed convex hull of the set $\left\{{0}\right\}\cup\left\{{\tfrac{t}{2^{n}}e_{n}}\;\middle|\;{n\in\mathbb{N}_{\geq 1}}\right\}$ in $\ell_{1}$. By discussion in [8, Section 2], $K_{t}$ is compact and has finite magnitude; specifically, by [8, Remark 2.2], we have $\mathsf{mag}_{\mathrm{Cmp}_{+}(\ell_{1})}(K_{t})\leq e^{t/2}$. Consider $K_{1}$; by Proposition 3.7, it is tractable, but is clearly not contained in a finite-dimensional vector subspace. Since $t\cdot K_{1}=K_{t}$ and $\lim_{t\to 0}e^{t/2}=1$, we see that nevertheless $K_{1}$ has the one-point property.
This raises the question of whether every non-empty compact subspace of a tractable subspace of $L_{1}$ (equivalently, by Proposition 3.4: every compact tractable subspace of $L_{1}$) has the one-point property. We do not yet know the answer, and the question is related to Conjecture 5.1 below. We do note, however, that the same idea as in the proof of [8, Theorem 3.1] does not work, as there it was crucial that the compact subspace was contained in a polytope (i.e. convex hull of finitely many points, and therefore necessarily contained in a finite-dimensional vector subspace). See, however, Corollary 1.4 in [14] for a partial result (cf. also Conjecture 1.5 in the same paper).
To wrap up this section, we note that in practice, it is not equally difficult to (dis)prove (uniform) continuity of magnitude for finite subspaces and compact subspaces. Specifically, it is usually easier to prove continuity for finite subspaces, and easier to disprove continuity for compact subspaces. For tractable spaces, the following proposition gives us the connection between the two, potentially allowing us to derive the more difficult case from the easier one. We will use this in the next section.
@@
    Conversely, suppose that $\mathsf{mag}_{\mathrm{Fin}_{+}(M)}\colon\mathrm{Fin}_{+}(M)\to\mathbb{R}$ is uniformly continuous/$C$-Lipschitz. Then it has a unique continuous extension $\mathrm{Cmp}_{+}(M)\to\mathbb{R}$ which is also uniformly continuous/$C$-Lipschitz. But for positive definite spaces this extension has to be the magnitude $\mathsf{mag}_{\mathrm{Cmp}_{+}(M)}\colon\mathrm{Cmp}_{+}(M)\to\mathbb{R}$ [12, Corollary 2.7].
@@ 4 Continuity of Magnitude on $\mathbb{R}$ and subspaces of $\mathbb{R}$
The goal of this section is to deduce that magnitude $\mathsf{mag}_{\mathrm{Cmp}_{+}(\mathbb{R})}\colon\mathrm{Cmp}_{+}(\mathbb{R})\to\mathbb{R}$ is continuous from the theory developed in the previous section. This is difficult to prove directly, even though the result has appeared in the literature previously (for example, see [6, Corollary 5.3.13] and [12, Corollary 2.12]). We also prove that the restrictions of $\mathsf{mag}_{\mathrm{Cmp}_{+}(\mathbb{R})}$ to bounded subspaces are Lipschitz.
@@
The space $\mathbb{R}_{\geq 0}$ is tractable, but the magnitude maps
$\mathsf{mag}_{\mathrm{Fin}_{+}(\mathbb{R}_{\geq 0})}\colon\mathrm{Fin}_{+}(\mathbb{R}_{\geq 0})\to\mathbb{R}$
and
$\mathsf{mag}_{\mathrm{Cmp}_{+}(\mathbb{R}_{\geq 0})}\colon\mathrm{Cmp}_{+}(\mathbb{R}_{\geq 0})\to\mathbb{R}$
are not uniformly continuous (therefore also not Lipschitz continuous). Hence, the same is also true for every tractable metric space containing an isometric copy of $\mathbb{R}_{\geq 0}$.
@@
$\big(1+2N\tanh(\tfrac{1}{2})\big)-\big(1+N\tanh(1)\big)=N\cdot\big(2\tanh(\tfrac{1}{2})-\tanh(1)\big)>C.$
@@
For all $\in\mathbb{R}, a<ba<b$, the magnitude maps $\mathsf{mag}_{\mathrm{Fin}_{+}({\mathbb{R}}_{[a,b]})}\colon\mathrm{Fin}_{+}({\mathbb{R}}_{[a,b]})\to\mathbb{R}$ and $\mathsf{mag}_{\mathrm{Cmp}_{+}({\mathbb{R}}_{[a,b]})}\colon\mathrm{Cmp}_{+}({\mathbb{R}}_{[a,b]})\to\mathbb{R}$ are Lipschitz with coefficient $1+\tfrac{b-a}{2}$.
@@
$r\cdot\Big(n-\sum_{i\in I}\frac{\tanh(\tfrac{a_{i+1}-a_{i}}{2}-r)-\tanh(\tfrac{a_{i+1}-a_{i}}{2})}{-r}-\sum_{j\in J}\big(1-\tfrac{1}{r}\big(\tfrac{a_{j+1}-a_{j}}{2}-\tanh(\tfrac{a_{j+1}-a_{j}}{2})\big)\big)\Big).$
@@
$1+\sum_{k\in{\mathbb{N}}_{[1,n)}}\tanh^{2}(\tfrac{a_{k+1}-a_{k}}{2})\leq 1+\sum_{k\in{\mathbb{N}}_{[1,n)}}\tanh(\tfrac{a_{k+1}-a_{k}}{2})\leq$
@@ 5 Conclusion
Although we have shown here continuity of magnitude only for the subspaces of $\mathbb{R}$, we suspect that it holds much more generally. We state this as the following conjecture, which is a generalization of [4, Conjecture 1.3] (see also discussions in [13], [14]).